# Stable dark matter from Pauli blocking in the degenerate fermion background with Quantum Field Theory

### v1

@@ Abstract
EPHOU-24-008
@@ 1 Introduction
Another possibility is to use the Pauli blocking in the DM decay to fermions [7, 8]. Bjaelde and Das studied the decay of DM into the Fermi sea of neutrinos, where the Pauli blocking controls the DM decay [7] which is analogous to the situtation of fermion production by preheating [9, 10]. They considered Yukawa couplings of the order of the unity and applied the change of the equation of state of DM at the last stage to find solutions of the cosmological problems. Ghosh and Mukhopadhyay studied DM momentum distribution from the decay of an inflaton, where the Pauli blocking modifies the momentum distribution of DM and its energy density [11]. Pauli blocking also affects the pair creation of fermion fields in strong electric field [12].
@@ 2 Particle Production with Boltzmann Equation
We consider a real scalar $\phi$ which can decay only into fermion pairs $\psi$ and $\bar{\psi}$. As a simple example, we assume the following Yukawa interaction with the coupling $\lambda$ in Lagrangian  [^1] We do not consider scattering processes, since those are subdominant for a very small Yukawa coupling as we consider in this paper.
$\begin{split}{\mathcal{L}}\ni\lambda\bar{\psi}\psi\phi.\end{split}$ (2.1)
@@
$\begin{split}\phi(p)\rightarrow\psi(k)+\bar{\psi}(k^{\prime}),\end{split}$ (2.2)
where $p=(m_{\phi},{\bf 0})$, $k=(E,\vec{\bf k})$, and $k^{\prime}=(E,-\vec{\bf k})$, with $|\vec{\bf k}|=p_{\psi}\equiv\frac{m_{\phi}}{2}\sqrt{1-\frac{4m_{\psi}^{2}}{m_{\phi}^{2}}}$ and $E^{2}=m_{\psi}^{2}+|\vec{\bf k}|^{2}$. The decay rate in the vacuum is given by
$\begin{split}\Gamma_{\phi,0}=\frac{p_{f}}{8\pi m_{\phi}^{2}}|\mathcal{M}|^{2}=\frac{\lambda^{2}m_{\phi}}{8\pi}\left(1-\frac{4m_{\psi}^{2}}{m_{\phi}^{2}}\right)^{3/2},\end{split}$ (2.3)
@@
$\begin{split}\tau_{\phi,0}=\frac{1}{\Gamma_{\phi,0}}\simeq 1.6\times 10^{13}\sec\left(\frac{10^{-12}}{\lambda}\right)^{2}\left(\frac{10^{-3}\,{\rm eV}}{m_{\phi}}\right).\end{split}$ (2.4)
@@
$\begin{split}\frac{\partial f_{\phi}(\vec{\bf p},t)}{\partial t}-Hp\frac{\partial f_{\phi}(\vec{\bf p},t)}{\partial p}=f_{\phi}^{\rm coll}(\vec{\bf p},t),\end{split}$ (2.5)
where the right-hand side is the collision term for $\phi$. For the process $\phi(\vec{\bf p})\leftrightarrow\psi(\vec{\bf k})+\bar{\psi}(\vec{\bf k^{\prime}})$, the collision term becomes
$\begin{split}f_{\phi}^{\rm coll}(\vec{\bf p},t)&=\frac{1}{2E_{\vec{\bf p}}}\int\frac{d^{3}\vec{\bf k^{\prime}}}{(2\pi)^{3}2E_{\vec{\bf k^{\prime}}}}\frac{d^{3}\vec{\bf k}}{(2\pi)^{3}2E_{\vec{\bf k}}}(2\pi)^{4}\delta^{4}(p-k^{\prime}-k)|\mathcal{M}|^{2},\\ &\times\left[f_{\psi}(\vec{\bf k},t)f_{\bar{\psi}}(\vec{\bf k}^{\prime},t)(1+f_{\phi}(\vec{\bf p},t))-f_{\phi}(\vec{\bf p},t)(1-f_{\psi}(\vec{\bf k},t))(1-f_{\bar{\psi}}(\vec{\bf k^{\prime}},t))\right],\end{split}$ (2.6)
@@
$\begin{split}|\mathcal{M}|^{2}=2\lambda^{2}m_{\phi}^{2}\left(1-\frac{4m_{\psi}^{2}}{m_{\phi}^{2}}\right).\end{split}$ (2.7)
@@
$\begin{split}n_{\phi}(t)=\int\frac{d^{3}\vec{\bf p}}{(2\pi)^{3}}f_{\phi}(\vec{\bf p},t).\end{split}$ (2.8)
In the case when $f_{\phi}(\vec{\bf p},t)\gg 1$, and $\phi$ is at rest, we can integrate the Boltzmann equation Eq. (2.5) to obtain its evolution equation
$\begin{split}\dot{n}_{\phi}(t)+3Hn_{\phi}(t)&\simeq-\Gamma_{\psi}n_{\phi}(t)\left.\left[1-2f_{\psi}(\vec{\bf k},t)\right]\right|_{|\vec{\bf k}|=p_{\psi}},\end{split}$ (2.9)
@@
$\begin{split}\frac{\partial f_{\psi}(\vec{\bf k},t)}{\partial t}-Hk\frac{\partial f_{\psi}(\vec{\bf k},t)}{\partial k}=\frac{1}{2}f_{\psi}^{\rm coll}(\vec{\bf k},t).\end{split}$ (2.10)
@@
$\begin{split}f_{\psi}^{\rm coll}(\vec{\bf k},t)&=-\frac{1}{2E_{\vec{\bf k}}}\int\frac{d^{3}\vec{\bf k^{\prime}}}{(2\pi)^{3}2E_{\vec{\bf k^{\prime}}}}\frac{d^{3}\vec{\bf p}}{(2\pi)^{3}2E_{\vec{\bf p}}}(2\pi)^{4}\delta^{4}(p-k^{\prime}-k)|\mathcal{M}|^{2}\\ &\times\left[f_{\psi}(\vec{\bf k},t)f_{\bar{\psi}}(\vec{\bf k^{\prime}},t)(1+f_{\phi}(\vec{\bf p},t))-f_{\phi}(\vec{\bf p},t)(1-f_{\psi}(\vec{\bf k},t))(1-f_{\bar{\psi}}(\vec{\bf k^{\prime}},t))\right],\end{split}$ (2.11)
@@
$\begin{split}f_{\psi}^{\rm coll}(\vec{\bf k},t)&\simeq\Gamma_{\psi}n_{\phi}(t)\left(\frac{2\pi^{2}}{p_{\psi}^{2}}\right)\delta(|\vec{\bf k}|-p_{\psi})\left[(1-f_{\psi}(\vec{\bf k},t))(1-f_{\bar{\psi}}(-\vec{\bf k},t))-f_{\psi}(\vec{\bf k},t)f_{\bar{\psi}}(-\vec{\bf k},t)\right],\\ &=\Gamma_{\psi}n_{\phi}(t)\left(\frac{2\pi^{2}}{p_{\psi}^{2}}\right)\delta(|\vec{\bf k}|-p_{\psi})\left[1-2f_{\psi}(\vec{\bf k},t)\right],\end{split}$ (2.12)
where $p_{\psi}$ is defined as
$\begin{split}p_{\psi}\equiv\frac{m_{\phi}}{2}\sqrt{1-\frac{4m_{\psi}^{2}}{m_{\phi}^{2}}}.\end{split}$ (2.13)
@@ 2.1 Without Cosmic Expansion
$\frac{\partial f_{\psi}(\vec{\bf k},t)}{\partial t}=\Gamma_{\psi}n_{\phi}(t)\left(\frac{\pi^{2}}{p_{\psi}^{2}}\right)\frac{1}{p_{\psi}}\delta(|\vec{\bf k}|/p_{\psi}-1)\left[1-2f_{\psi}(\vec{\bf k},t)\right],$ (2.14)
@@
$\begin{split}\dot{n}_{\phi}(t)&\simeq-\Gamma_{\psi}n_{\phi}(t)\left.\left[1-2f_{\psi}(\vec{\bf k},t)\right]\right|_{|\vec{\bf k}|=p_{\psi}}.\end{split}$ (2.15)
If we integrate the sum of Eq. (2.15) and Eq. (2.14) times $\frac{p_{\psi}^{2}}{\pi^{2}}$, we obtain the equation of total number density conservation given by
$\begin{split}n_{\phi}(t)+\frac{p_{\psi}^{3}}{\pi^{2}}\left.f_{\psi}(\vec{\bf k},t)\right|_{|\vec{\bf k}|=p_{\psi}}=n_{\phi}(t_{i})+\frac{p_{\psi}^{3}}{\pi^{2}}\left.f_{\psi}(\vec{\bf k},t_{i})\right|_{|\vec{\bf k}|=p_{\psi}}\equiv C_{0}.\end{split}$ (2.16)
Using Eq. (2.16), we can replace $f_{\psi}$ in Eq. (2.15) by $n_{\phi}$ to obtain an differential equation for $n_{\phi}(t)$
$\begin{split}\dot{n}_{\phi}(t)&\simeq-\Gamma_{\psi}n_{\phi}(t)\left.\left[1-2\{C_{0}-n_{\phi}(t)\}\frac{\pi^{2}}{p_{\psi}^{3}}\right]\right|_{|\vec{\bf k}|=p_{\psi}}\\ &=-An_{\phi}(t)+Bn_{\phi}^{2}(t),\end{split}$ (2.17)
@@
$A\equiv\Gamma_{\psi}\left(1-C_{0}\frac{2\pi^{2}}{p_{\psi}^{3}}\right),\qquad\mathrm{and}\qquad B\equiv-\Gamma_{\psi}\frac{2\pi^{2}}{p_{\psi}^{3}}.$ (2.18)
@@
$\begin{split}\frac{1}{n_{\phi}(t)}&=\frac{B}{A}+\left(\frac{1}{n_{\phi}(t_{i})}-\frac{B}{A}\right)\exp[A(t-t_{i})]\\ =&\left[C_{0}-\frac{p_{\psi}^{3}}{2\pi^{2}}\right]^{-1}+\left(\frac{1}{n_{\phi}(t_{i})}-\left[C_{0}-\frac{p_{\psi}^{3}}{2\pi^{2}}\right]^{-1}\right)\exp\left[\Gamma_{\psi}\left(1-C_{0}\frac{2\pi^{2}}{p_{\psi}^{3}}\right)(t-t_{i})\right],\end{split}$ (2.19)
@@
$\begin{split}(I)\quad C_{0}<\frac{p_{\psi}^{3}}{2\pi^{2}}:&\quad n_{\phi}(t)\rightarrow 0,\quad f_{\psi}\rightarrow\frac{\pi^{2}}{p_{\psi}^{3}}C_{0},\\ &\quad n_{\psi}(t)\rightarrow g\int\frac{d^{3}k}{(2\pi)^{3}}\frac{\pi^{2}}{p_{\psi}^{3}}C_{0}\delta(|\vec{\bf k}|/p_{\psi}-1)=C_{0},\\ (II)\quad C_{0}>\frac{p_{\psi}^{3}}{2\pi^{2}}:&\quad n_{\phi}(t)\rightarrow C_{0}-\frac{p_{\psi}^{3}}{2\pi^{2}}=n_{\phi}(t_{i})-\frac{p_{\psi}^{3}}{\pi^{2}}\left(\frac{1}{2}-\left.f_{\psi}(\vec{\bf k},t_{i})\right|_{|\vec{\bf k}|=p_{\psi}}\right),\\ &\quad f_{\psi}\rightarrow\frac{1}{2},\quad n_{\psi}(t)\rightarrow g\int\frac{d^{3}k}{(2\pi)^{3}}\frac{1}{2}\delta(|\vec{\bf k}|/p_{\psi}-1)=\frac{p_{\psi}^{3}}{2\pi^{2}},\end{split}$ (2.20)
@@
In Fig. 1, we show the evolution of the number density for $\phi$ and $\psi$ normalized by $p_{\psi}^{3}/2\pi^{2}$, for both case with (I) $C_{0}<p_{\psi}^{3}/2\pi^{2}$ (orange) and (II) $C_{0}>p_{\psi}^{3}/2\pi^{2}$ (blue). For case (I) $\phi$ completely decays to fermion, while for case (II) only small fraction of $\phi$ decays to fermion whose phase space is fully occupied.
@@
$\begin{split}n_{\psi}(t)&=n_{\bar{\psi}}(t)\simeq\frac{p_{\psi}^{3}}{2\pi^{2}}\left(1-\exp[-n_{\phi}(t_{i})\Gamma_{\phi}\frac{\pi^{2}}{p_{\psi}^{3}}(t-t_{i})]\right),\\ n_{\phi}(t)&\simeq n_{\phi}(t_{i})-n_{\psi}(t).\end{split}$ (2.21)
Initially, $n_{\psi}$ grows linearly with time as $n_{\psi}\sim n_{\phi}(t_{i})\Gamma_{\phi}t$ and then later fills the Fermi surface nearly at the time scale
$\begin{split}t_{c}\simeq\left(\frac{n_{\phi}(t_{i})}{p_{\psi}^{3}/\pi^{2}}\right)^{-1}\Gamma_{\phi}^{-1}\gg\Gamma_{\phi}^{-1},\end{split}$ (2.22)
compared to $\Gamma_{\phi}^{-1}$ when there is no Pauli blocking.

### v2

@@ Abstract
keywords:
Dark matter , Pauli blocking , Boltzmann equation , Quantum field theory , Neutrino
journal: Physics of the Dark Universe
[…]
@@ 1 Introduction
Another possibility is to use the Pauli blocking in the DM decay to fermions [7, 8]. Bjaelde and Das studied the decay of DM into the Fermi sea of neutrinos, where the Pauli blocking controls the DM decay [7] which is analogous to the situation of fermion production by preheating [9, 10]. They considered Yukawa couplings of the order of the unity and applied the change of the equation of state of DM at the last stage to find solutions of the cosmological problems. Ghosh and Mukhopadhyay studied DM momentum distribution from the decay of an inflaton, where the Pauli blocking modifies the momentum distribution of DM and its energy density [11]. Pauli blocking also affects the pair creation of fermion fields in strong electric field [12].
@@ 2 Particle Production with Boltzmann Equation
We consider a real scalar $\phi$ which can decay only into fermion pairs $\psi$ and $\bar{\psi}$. As a simple example, we assume the following Yukawa interaction with the coupling $\lambda$ in Lagrangian  [^1] We do not consider scattering processes, since those are subdominant for small Yukawa coupling as we consider in this paper.
$\begin{split}{\mathcal{L}}\ni\lambda\bar{\psi}\psi\phi.\end{split}$ (1)
@@
$\begin{split}\phi(p)\rightarrow\psi(k)+\bar{\psi}(k^{\prime}),\end{split}$ (2)
where $p=(m_{\phi},{\bf 0})$, $k=(E,\vec{\bf k})$, and $k^{\prime}=(E,-\vec{\bf k})$, with $|\vec{\bf k}|=p_{\psi}\equiv\frac{m_{\phi}}{2}\sqrt{1-R^{2}}$, $E^{2}=m_{\psi}^{2}+|\vec{\bf k}|^{2}$, and $R\equiv 2m_{\psi}/m_{\phi}$. The decay rate in the vacuum is given by
$\begin{split}\Gamma_{\phi}=\frac{p_{\psi}}{8\pi m_{\phi}^{2}}|\mathcal{M}|^{2}=\frac{\lambda^{2}m_{\phi}}{8\pi}\left(1-R^{2}\right)^{3/2},\end{split}$ (3)
@@
$\begin{split}\tau_{\phi,0}=\frac{1}{\Gamma_{\phi}}\simeq 1.6\times 10^{13}\sec\left(\frac{10^{-12}}{\lambda}\right)^{2}\left(\frac{10^{-3}\,{\rm eV}}{m_{\phi}}\right)\left(1-R^{2}\right)^{-3/2}.\end{split}$ (4)
@@
$\begin{split}\frac{\partial f_{\phi}(\vec{\bf p},t)}{\partial t}-Hp\frac{\partial f_{\phi}(\vec{\bf p},t)}{\partial p}=f_{\phi}^{\rm coll}(\vec{\bf p},t),\end{split}$ (5)
where the right-hand side is the collision term for $\phi$ and $H$ is the cosmic expansion rate. For the process $\phi(\vec{\bf p})\leftrightarrow\psi(\vec{\bf k})+\bar{\psi}(\vec{\bf k^{\prime}})$, the collision term becomes
$\begin{split}f_{\phi}^{\rm coll}(\vec{\bf p},t)&=\frac{1}{2E_{\vec{\bf p}}}\int\frac{d^{3}\vec{\bf k^{\prime}}}{(2\pi)^{3}2E_{\vec{\bf k^{\prime}}}}\frac{d^{3}\vec{\bf k}}{(2\pi)^{3}2E_{\vec{\bf k}}}(2\pi)^{4}\delta^{4}(p-k^{\prime}-k)|\mathcal{M}|^{2}\\ &\times\left[f_{\psi}(\vec{\bf k},t)f_{\bar{\psi}}(\vec{\bf k}^{\prime},t)(1+f_{\phi}(\vec{\bf p},t))-f_{\phi}(\vec{\bf p},t)(1-f_{\psi}(\vec{\bf k},t))(1-f_{\bar{\psi}}(\vec{\bf k^{\prime}},t))\right],\end{split}$ (6)
@@
$\begin{split}|\mathcal{M}|^{2}=2\lambda^{2}m_{\phi}^{2}\left(1-R^{2}\right).\end{split}$ (7)
@@
$\begin{split}n_{\phi}(t)=\int\frac{d^{3}\vec{\bf p}}{(2\pi)^{3}}f_{\phi}(\vec{\bf p},t).\end{split}$ (8)
In the case when $f_{\phi}(\vec{\bf p},t)\gg 1$, and $\phi$ is at rest, by integrating the Boltzmann Eq. (5), we obtain its evolution equation
$\begin{split}\dot{n}_{\phi}(t)+3Hn_{\phi}(t)&\simeq-\Gamma_{\phi}n_{\phi}(t)\left.\left[1-2f_{\psi}(\vec{\bf k},t)\right]\right|_{|\vec{\bf k}|=p_{\psi}},\end{split}$ (9)
@@
$\begin{split}\frac{\partial f_{\psi}(\vec{\bf k},t)}{\partial t}-Hk\frac{\partial f_{\psi}(\vec{\bf k},t)}{\partial k}=\frac{1}{2}f_{\psi}^{\rm coll}(\vec{\bf k},t).\end{split}$ (10)
@@
$\begin{split}f_{\psi}^{\rm coll}(\vec{\bf k},t)&=-\frac{1}{2E_{\vec{\bf k}}}\int\frac{d^{3}\vec{\bf k^{\prime}}}{(2\pi)^{3}2E_{\vec{\bf k^{\prime}}}}\frac{d^{3}\vec{\bf p}}{(2\pi)^{3}2E_{\vec{\bf p}}}(2\pi)^{4}\delta^{4}(p-k^{\prime}-k)|\mathcal{M}|^{2}\\ &\times\left[f_{\psi}(\vec{\bf k},t)f_{\bar{\psi}}(\vec{\bf k^{\prime}},t)(1+f_{\phi}(\vec{\bf p},t))-f_{\phi}(\vec{\bf p},t)(1-f_{\psi}(\vec{\bf k},t))(1-f_{\bar{\psi}}(\vec{\bf k^{\prime}},t))\right],\end{split}$ (11)
@@
$\begin{split}f_{\psi}^{\rm coll}(\vec{\bf k},t)&\simeq\Gamma_{\psi}n_{\phi}(t)\left(\frac{2\pi^{2}}{p_{\psi}^{2}}\right)\delta(|\vec{\bf k}|-p_{\psi})\left[(1-f_{\psi}(\vec{\bf k},t))(1-f_{\bar{\psi}}(-\vec{\bf k},t))-f_{\psi}(\vec{\bf k},t)f_{\bar{\psi}}(-\vec{\bf k},t)\right]\\ &=\Gamma_{\psi}n_{\phi}(t)\left(\frac{2\pi^{2}}{p_{\psi}^{2}}\right)\delta(|\vec{\bf k}|-p_{\psi})\left[1-2f_{\psi}(\vec{\bf k},t)\right].\end{split}$ (12)
@@ 2.1 Without Cosmic Expansion
$\frac{\partial f_{\psi}(\vec{\bf k},t)}{\partial t}=\Gamma_{\phi}n_{\phi}(t)\left(\frac{\pi^{2}}{p_{\psi}^{2}}\right)\frac{1}{p_{\psi}}\delta(|\vec{\bf k}|/p_{\psi}-1)\left[1-2f_{\psi}(\vec{\bf k},t)\right],$ (13)
@@
$\begin{split}\dot{n}_{\phi}(t)&\simeq-\Gamma_{\psi}n_{\phi}(t)\left.\left[1-2f_{\psi}(\vec{\bf k},t)\right]\right|_{|\vec{\bf k}|=p_{\psi}}.\end{split}$ (14)
If we integrate the sum of Eq. (14) and Eq. (13) times $\frac{p_{\psi}^{3}}{\pi^{2}}$, we obtain the equation of total number density conservation given by
$\begin{split}n_{\phi}(t)+\frac{p_{\psi}^{3}}{\pi^{2}}\left.f_{\psi}(\vec{\bf k},t)\right|_{|\vec{\bf k}|=p_{\psi}}=n_{\phi}(t_{i})+\frac{p_{\psi}^{3}}{\pi^{2}}\left.f_{\psi}(\vec{\bf k},t_{i})\right|_{|\vec{\bf k}|=p_{\psi}}\equiv C_{0}.\end{split}$ (15)
Using Eq. (15), we can replace $f_{\psi}$ in Eq. (14) by $n_{\phi}$ to obtain an differential equation for $n_{\phi}(t)$
$\begin{split}\dot{n}_{\phi}(t)&\simeq-\Gamma_{\psi}n_{\phi}(t)\left.\left[1-2\{C_{0}-n_{\phi}(t)\}\frac{\pi^{2}}{p_{\psi}^{3}}\right]\right|_{|\vec{\bf k}|=p_{\psi}}\\ &=-An_{\phi}(t)+Bn_{\phi}^{2}(t),\end{split}$ (16)
@@
$A\equiv\Gamma_{\psi}\left(1-C_{0}\frac{2\pi^{2}}{p_{\psi}^{3}}\right),\qquad\mathrm{and}\qquad B\equiv-\Gamma_{\psi}\frac{2\pi^{2}}{p_{\psi}^{3}}.$ (17)
@@
$\begin{split}\frac{1}{n_{\phi}(t)}&=\frac{B}{A}+\left(\frac{1}{n_{\phi}(t_{i})}-\frac{B}{A}\right)\exp[A(t-t_{i})]\\ =&\left[C_{0}-\frac{p_{\psi}^{3}}{2\pi^{2}}\right]^{-1}+\left(\frac{1}{n_{\phi}(t_{i})}-\left[C_{0}-\frac{p_{\psi}^{3}}{2\pi^{2}}\right]^{-1}\right)\exp\left[\Gamma_{\psi}\left(1-C_{0}\frac{2\pi^{2}}{p_{\psi}^{3}}\right)(t-t_{i})\right],\end{split}$ (18)
@@
$\begin{split}(I)\quad C_{0}<\frac{p_{\psi}^{3}}{2\pi^{2}}:&\quad n_{\phi}(t)\rightarrow 0,\quad f_{\psi}\rightarrow\frac{\pi^{2}}{p_{\psi}^{3}}C_{0},\\ &\quad n_{\psi}(t)\rightarrow g\int\frac{d^{3}k}{(2\pi)^{3}}\frac{\pi^{2}}{p_{\psi}^{3}}C_{0}\delta(|\vec{\bf k}|/p_{\psi}-1)=C_{0},\\ (II)\quad C_{0}>\frac{p_{\psi}^{3}}{2\pi^{2}}:&\quad n_{\phi}(t)\rightarrow C_{0}-\frac{p_{\psi}^{3}}{2\pi^{2}}=n_{\phi}(t_{i})-\frac{p_{\psi}^{3}}{\pi^{2}}\left(\frac{1}{2}-\left.f_{\psi}(\vec{\bf k},t_{i})\right|_{|\vec{\bf k}|=p_{\psi}}\right),\\ &\quad f_{\psi}\rightarrow\frac{1}{2},\quad n_{\psi}(t)\rightarrow g\int\frac{d^{3}k}{(2\pi)^{3}}\frac{1}{2}\delta(|\vec{\bf k}|/p_{\psi}-1)=\frac{p_{\psi}^{3}}{2\pi^{2}},\end{split}$ (19)
@@
In Fig. 1, we show the evolution of the number density for $\phi$ and $\psi$ normalized by $p_{\psi}^{3}/2\pi^{2}$, for both case with (I) $C_{0}<p_{\psi}^{3}/2\pi^{2}$ (orange) and (II) $C_{0}>p_{\psi}^{3}/2\pi^{2}$ (blue). For case (I), $\phi$ completely decays to fermion, while for case (II) only small fraction of $\phi$ decays to fermion whose phase space is fully occupied.
@@
$\begin{split}n_{\psi}(t)&=n_{\bar{\psi}}(t)\simeq\frac{p_{\psi}^{3}}{2\pi^{2}}\left(1-\exp[-n_{\phi}(t_{i})\Gamma_{\phi}\frac{\pi^{2}}{p_{\psi}^{3}}(t-t_{i})]\right),\\ n_{\phi}(t)&\simeq n_{\phi}(t_{i})-n_{\psi}(t).\end{split}$ (20)
Initially, $n_{\psi}$ grows linearly with time as $n_{\psi}\sim n_{\phi}(t_{i})\Gamma_{\phi}t$ and later fills the Fermi surface nearly at the time scale
$\begin{split}t_{c}\simeq\left(\frac{n_{\phi}(t_{i})}{p_{\psi}^{3}/\pi^{2}}\right)^{-1}\Gamma_{\phi}^{-1},\end{split}$ (21)
which is smaller than the lifetime of scalar when there is no Pauli blocking, $\Gamma_{\phi}^{-1}$.
Note that the scatterings $\phi\phi\leftrightarrow\psi\psi$ and $\phi\psi\leftrightarrow\phi\psi$ can happen in the order of $\lambda^{4}$ and does not affect our results once the timescale of the scattering is much smaller than the Hubble time. However, in the opposite case when the scattering time-scale becomes shorter than the expansion, the scatterings can redistribute the momentum of $\phi$ and $\psi$, which can initiate the decrease of condensate $\phi$ [8].